OpenStudy (lilai3):

Can someone please help me...please? How many ways can 6 students be arranged in a lunch line? A family has a bike rack that fits 7 bikes but they only have 5 bikes. How many ways can the bikes fit in the bike rack? How many 4-digit pin numbers are available if no number is repeated?

OpenStudy (lilai3):

helllpp mee please!!!!!!

OpenStudy (anonymous):

lunch line: 6! = 720

OpenStudy (lilai3):

thhaaaannnnkkk yoooouuuuu ,and also the other problems... and could you possibly explain to me plz? (i'll give u a meeedddaaalll?????????) >:D

OpenStudy (anonymous):

it's the formula such that having x to arranged, then we have x factorial (x!) so if it's 10 students then 10!. 20 will be 20!, and so on

OpenStudy (anonymous):

pin number we have 10 digits number 0-9 so we can choose the first number from 10 numbers then we choose the 2nd number from 9 numbers (since no number repeated) then we choose the 3rd number from 8 numbers then we choose the last number from 7 numbers left finally we have: 10*9*8*7 = 5040

OpenStudy (anonymous):

we the numbers is allowed to be repeated, then we have 10 ways to choose each number, and so: 10*10*10*10 = 10^4 = 10000

OpenStudy (lilai3):

wait... r u talking about the bike one?

OpenStudy (anonymous):

no the pin number

OpenStudy (lilai3):

oh... so then the real answer's suppose to be 10000... rite?

OpenStudy (anonymous):

no the answer for your pin number is 5040 the 10000 is the insight idea

OpenStudy (lilai3):

oh i get it now.. so If the numbers can be repeated, dat means dat dere are 10,000 possiblities, correct?

OpenStudy (anonymous):

yes

OpenStudy (lilai3):

so then wut about the biking one?

OpenStudy (anonymous):

do you know about permutation?

OpenStudy (lilai3):

ya

OpenStudy (lilai3):

is it 7*6*5*4*3=2520?

OpenStudy (anonymous):

yes, the same result if you apply 7P5 = 7!/(7-5)!

OpenStudy (lilai3):

oh, okay then and thank you but i also haf 2 more questions if u don't mind: How many ways can you select 3 sheriff deputies from 8 candidates? How many handshakes can occur between 5 people if everyone shakes hands? And can u also explain each one? i would be grateful and thank you! :D

OpenStudy (anonymous):

deputies candidate: having 3 choosing from 8 so it's 8C3 = 8!/[3!*(8-3)! = 112

OpenStudy (lilai3):

i'm sorry, i can't understand all dat... stuff

OpenStudy (anonymous):

you need to choose 3 people from 8 candidates and 8C3 is the formula, the same formula with permutation

OpenStudy (lilai3):

how did u get 112?

OpenStudy (anonymous):

8C3 = 8!/[3!*(8-3)! = 8!/[3!*5! = (8*7*6*5*4*3*2*1) / [(3*2*1)*(5*4*3*2*1)] = 8*7*6 / 3*2 = 8*7 = 56, this is the correct answer

OpenStudy (lilai3):

oh thank you and good night and then i'll be off.. with my personal stuff....... i'll sorta "contact" you if I'll need anymore help... ;)

OpenStudy (anonymous):

ok, good night

OpenStudy (lilai3):

waaaaaiiiiitttttt...... you forgot to explain the problem about the handshakes!!!!!

OpenStudy (lilai3):

is it 3125? because then it's like 5*5*5*5*5...right?

OpenStudy (anonymous):

we need to have 2 people to shake hand. So we need to choose 2 people from 5 people to have the hand shake. so 5C2 = 5!/[2!*(5-2)!] = 5!/(2!*3!) = 5*4*3*2*1 / [(2*1)*(3*2*1)] = 5*4 / 2 = 10

OpenStudy (lilai3):

oh never mind i get it then thanks... so its: (5*4)/(2*1)=10 rite?

OpenStudy (lilai3):

no answer? fne good night then